# Euclid Quick Data Release (Q1)

### v2

@@ 4.2 Comparisons with other classifications
In this work, when referring to completeness against a specific reference sample (e.g., a catalog of selected AGN candidates), we mean the fraction of objects in that reference sample that are successfully identified by our method. This is due to the true astrophysical completeness being unknown and extremely difficult to measure from the observational data alone. Our context is equivalent to the standard machine learning metric of recall for the positive class, which is defined in Eq. 13.
When developing a method for initial candidate selection, where the goal is to flag objects for follow-up or multi-wavelength analysis, high recall is often prioritised. The aim is to minimise the loss of potentially interesting or rare sources (i.e., reduce false negatives), even if this comes at the cost of lower precision (i.e., including a larger fraction of contaminants or false positives in the selected sample). These contaminants can often be filtered out through subsequent analysis stages or by incorporating additional data. Our primary goal here is to demonstrate the method’s capability to recover a wide variety of known AGN and QSO types, making recall a key indicator of its potential utility in broad searches.
As is particularly true for AGN and QSO, it is impossible to create a selection criterion that will identify a single population of objects with high completeness and high purity. Instead, we will use the aforementioned samples, many of which are only candidates themselves, to assess the contamination and purity of our method to the extent that is possible.
@@ 4.3 Data expectations
As a baseline for our morphology-based approach, it is first necessary to verify that the Euclid VIS images contain sufficient information to resolve the contrast between the compact central emission characteristic of an AGN and the broader light profiles of typical galaxies. We analyse the relationship between the brightest central pixel and its surrounding pixels to quantify a difference between normal galaxies and AGN. By producing ratios comparing the central pixel to 1 and 2-pixels-wide regions around the central pixel, we form a rate of change (ROC) metric over sources, as described by
@@
Pixels surrounding the brightest pixel, which share values similar to the maximum, will produce a ROC value closer to zero, implying that the source does not feature a prominent active component. Sources featuring a single very bright pixel surrounded by fainter pixels will have a ROC value closer to its maximum of 2. The produced distributions of sources, analysed according to both $I_{\scriptscriptstyle\rm E}$ magnitude and various selection criteria (Sect. 4.2), are shown in Fig. 3. As expected, sources with a high flux point-spread function (PSF) light profile show the highest ROC values. This effect is seen in the star sample and the QSO-dominated selections (R90 WISE, Quaia, DESI-QSO, and Euclid-QSO). On the other hand, spectroscopically confirmed normal galaxies show very broad ROC distributions.
@@ 4.4 Addressing the use of asinh for image scaling
Figure 4: Distribution differences of the rate of change value over each class. The narrow peaks from the non-galaxy class are significantly widened after applying the asinh transformation, causing a large overlap between classes.
Table 1: Percentage of galaxies that have rate of change value higher than the mean rate of change of the other classes. Due to the ability to compress the dynamic range of the image, the separability of the classes becomes more difficult.
@@
AGN 0.75% 1.15%
QSO 0.20% 7.82%
Stars 0.17% 1.86%
Given its prevalence in galaxy morphology-based research (Lupton et al. 2004), the effect of using an $\operatorname{asinh}$ transformation on the data is also investigated. The function is often used to benefit visual inspections by reducing the difference between the brightest and faintest pixels. In our approach, however, when attempting to differentiate galaxy and non-galaxy sources, the $\operatorname{asinh}$ function becomes a hindrance, causing the distributions to more significantly overlap, as shown in Fig. 4. To generate a measure for class overlap, we measure the ROC for each image, taking the mean of each non-galaxy class. We use the percentage of our true galaxy sample with a ROC value over the respective class mean as our overlap metric. The extent of the overlap is showcased in Table 1, where galaxy overlap increases by at least a factor of 2.5 and in the extreme case, nearly 50 times more galaxies overlap with stars.
In the following sections of the paper, we will discuss the results of the diffusion model and analyse the performance of the inpainting pipeline that automatically learns to reconstruct the central emission expected for the normal galaxy population, allowing us to identify stars and AGN as outliers.
@@ 5 Results
Although the focus of the paper is on the model’s performance on small areas of a galaxy’s core, Repaint’s adaptability allows the pipeline to handle a variety of mask shapes and sizes. More general examples of the inpainting performance can be seen in Fig. 17. With larger masks, fewer pixels remain for conditioning, making the output more susceptible to artifacts. However, Lugmayr et al. (2022) demonstrate that this can be mitigated by increasing the number of resamples during the Repaint loop, at the cost of longer inference times.
@@ 5.1 Inpainting of AGN
For our inpainting, we use a $\times5$ pixel mask centred on the brightest pixel, following the constraints described in Sect. 3.1. Comparing the expected pixel values with the generated outputs, Fig. 5 shows the model’s performance. The assumption that the model will predict a fainter core than the original does appear to be true, with a near-global bias for underpredicting, as shown by the dashed $y=x$ line. The majority of sources behave in a consistent way, with the exception of a smaller cluster of sources that predict more than an order of magnitude fainter than the original. This distinct cluster could provide sufficient candidates for AGN.
@@
To analyse how the error of each subset of class candidates have performed, comparisons are made between the original pixels and inpainted reconstruction for each image. The MSE of the pixel values are shown in Fig. 6. Each subset is binned according to its $I_{\scriptscriptstyle\rm E}$ magnitude, providing a distinction between the relative brightness of each image.
@@
There is a clear distribution difference between the respective class candidates. Brighter sources, particularly stars and luminous QSOs, produce higher errors, followed by fainter QSOs and the lowest-magnitude AGN candidates. Galaxies and the faintest AGN candidates show significantly lower errors, showing the model’s tendancy to reconstruct sources without strong AGN features more effectively. This indicates that the model has learned the anticipated bias introduced by the uneven distribution within the data set. While MSE is scale-dependant and influenced by the source’s VIS $I_{\scriptscriptstyle\rm E}$ magnitude, the separation between the errors from the galaxy class and those of the other classes suggests it remains useful as a selection metric. The presence of low-error AGN candidates may also indicate contamination within the AGN labels.
@@
It is acknowledged that because the model is trained on the full, unlabeled galaxy population, its prior distribution does contain information about AGN, even if they are rare. This could, in principle, lead the model to occasionally reconstruct AGN-like features, resulting in false negatives. However, the strong performance in separating known AGN and QSO populations, as shown in Figure 6, suggests that the prior is sufficiently dominated by non-active galaxies for this outlier detection framework to be effective.
@@
To showcase how the performance of the inpainting method impacts the overall image, Fig. 7 presents the input, output and residuals for a selection of sources, including a close-up of the masked area both before and after inpainting. It is clear how bright central cores can be entirely removed with inpainting and replaced by regions consistent with the surrounding pixels. The model can also accurately recreate images that do not feature any significant core. This shows that the model has a good internal representation of galaxies and how they should behave.
@@
Figure 9: Threshold produced between the top right cluster of points and the largely galaxy-dominated cluster. The coloured dots represent the VIS $I_{\scriptscriptstyle\rm E}$ magnitude for each positive candidate for the respective selections. The grey dots show the values of all the images in the data set. The majority of the sources within each non-galaxy class are captured with this simple linear boundary, allowing for high recall across all classes. The sources not captured are typically those of higher VIS magnitude and, therefore, fainter. The tail formed, where sources exhibit a maximum pixel difference ratio $10^{0}$, shows outlier sources with brighter predictions after inpainting.
@@ 5.2 Max pixel difference ratio
Although reconstruction error allows stars and a subset of AGN to stand out, a substantial overlap between suspected AGN and DESI-classified galaxies remains, especially at higher magnitudes. Even though measured on a pixel-by-pixel scale, MSE averages away specific pixel differences, leading to less nuanced differences between input and output. For example, if the reconstructed image ended up being the correct values but shifted by one pixel, this would be penalised heavily under MSE. From the perspective of this paper, such a scenario would represent a desirable outcome, as the model would correctly predict the presence of a bright, unresolved component, even if its exact location were shifted by a few pixels within the mask. Another example would be if only the brightest pixel is different in the reconstruction. MSE would significantly reduce the impact of this measurement, potentially scoring similarly to a source that gave minor errors in every pixel within the mask.
@@
To ensure such cases are not overlooked, we measure the pixel value ratio between the brightest pixel from the input mask and the brightest pixel from the reconstruction. As shown in Fig. 8, the produced distributions show a much tighter spread of values for many of the AGN selections, with a skew towards higher scores. The majority of the galaxy subset produces a ratio less than 10, whereas a large population of suspected AGN from each selection feature a larger ratio.
@@
Another important difference between this metric and MSE is how the results are relative, reducing the impact of magnitude on the recorded values. Many of the high-density peaks remain consistent, even as the magnitude of the sources increases. This implies that this metric is more robust at detecting fainter AGN than MSE.
@@
The distribution also highlights the scale of the extreme values, with many sources showing pixel values approaching 100 times greater than their counterparts. Exploring these large disparities can provide insights into the types of sources that are significantly sensitive to the model’s capabilities. Sources producing the highest values for the ratio are those where the reconstruction produces a cutout significantly fainter than the original image. As seen in Fig. 20, this space is dominated by point-like objects. Having the sharp falloff of pixel intensity in such a small area leads to the mask covering the majority of the source, if not the entirety. When the model attempts to reconstruct the missing pixels, it generates values that are coherent with the surrounding pixels. In these cases, the point-like source contributed very little to the surrounding pixels, leading to the generation of only background noise without any source.
@@
The behaviour of the previous examples matches the model’s expected behaviour and the paper’s initial assumptions. Applying inpainting to this task is expected to produce outputs that are similar to, or produce a flattened light profile of, the original pixels. However, a subset of sources exists that exhibit the opposite behaviour and generate a much brighter collection of pixels. Figure 21 shows examples of such sources. When analysing the sources that exhibit this behaviour, we found that they have overwhelmingly low S/N images. Further exploration of the impact of S/N on the diffusion process is shown in Sect. 6.1.
@@
Table 2: The number of sources that produced brighter pixels in the masked area after repainting. The normalised loss produces a clear advantage in reducing the number of erroneous results.
@@
The number of sources whose max pixel difference ratio is less than one for models trained on the two loss functions discussed in Sect. 2.3 are shown in Table 2. The normalised loss significantly reduces the number of sources exhibiting this behaviour. The normalised loss trained on the complete data set was selected as our applied model because it further reduced the number of erroneous predictions.
@@ 5.3 Morphology-based performance
Given the variety of shapes and sizes of objects that will be witnessed using Euclid, it is important to verify that the model can handle the complexities of different galaxy morphologies. Utilising the refined morphology classifications provided from Euclid Collaboration: Walmsley et al. (2025), we provide examples of edge-on (disk-edge-on_yes_fraction >0.5>0.5), spiral (has-spiral-arms_yes_fraction >0.7>0.7), and mergers (merging_merger_fraction >0.3>0.3) in Figs. 22, 23, LABEL:, and 24, respectively.
@@ 5.4 Creating a classifier
Due to the difference in distributions between the two metrics, any non-overlapping cases can be found when used in combination. As seen in Fig. 9, there is a cluster of sources in the top right. The key aspect of this cluster is its clear boundary from the DESI galaxy selection. This is vital for ensuring high reliability in classifications. We create a simple decision boundary between the two clusters of points, as shown in Fig. 9. Although a more fine-tuned boundary would be possible, especially when incorporating a secondary machine-learning model, it is important to show the benefits of utilising the inpainting method through only its error metrics. The equation of our linear boundary is y=0.01x+9y=-0.01x+9.
@@ 5.5 Variance of reconstructions
7 samples are generated as part of a single application of the pipeline. To ensure that any low variance in samples is not caused by the model simply ‘remembering’ the true image from its training examples, all resampling is applied on images the model has never seen before. The mean and variance over a normalised-by-brightness MSE are shown in Fig. 10.
The top figure shows that irrespective of magnitude, the model achieves similar relative median error across all images, highlighting the impact of the normalised loss throughout training. However, the difference in the scheduler over images of different quality and brightness is evident in the bottom figure. The consistency of generated outputs varies drastically as fainter (and therefore typically lower S/N) images produce widely different pixels when resampled.
Figure 10: Variations in the output error from repeat samples of the same fixed pixels. Over seven separate inpainting runs, the model shows consistent median error when normalised by each respective magnitude (top panel). The variance in the output errors is less consistent with the brighter, lower magnitudes, which provide more robust inpainting than those of fainter images.
Figure 11: Comparing the PSF total flux with the Srsic flux provides a clear separation between the galaxy class and the majority of the non-galaxy selection. This allows for a simple linear boundary for classifying potential AGNs, with the majority being point-like objects. However, it is most apparent with selections from AllWISE R90 and X-ray-selected AGN that many sources would not be selected using these features. The coloured dots represent the VIS $I_{\scriptscriptstyle\rm E}$ magnitude for each positive candidate for the respective selections. The grey dots show the scores of all the images in the data set.

### v3

@@ 4.2 Comparisons with other classifications
In this work, when referring to completeness against a specific reference sample (e.g. a catalogue of selected AGN candidates), we mean the fraction of objects in that reference sample that are successfully identified by our method. This is due to the true astrophysical completeness being unknown and extremely difficult to measure from the observational data alone. Our context is equivalent to the standard machine learning metric of recall for the positive class, which is defined in Eq. 13.
When developing a method for initial candidate selection, whereby the goal is to flag objects for follow-up or multi-wavelength analysis, high recall is often prioritised. The aim is to minimise the loss of potentially interesting or rare sources (i.e. reduce false negatives), even if this comes at the cost of lower precision (i.e. including a larger fraction of contaminants or false positives in the selected sample). These contaminants can often be filtered out through subsequent analysis stages or by incorporating additional data. Our primary goal here is to demonstrate the method’s capability to recover a wide variety of known AGN and QSO types, making recall a key indicator of its potential utility in broad searches.
As is particularly true for AGN and QSOs, it is impossible to create a selection criterion that will identify a single population of objects with high completeness and high purity. Instead, we shall use the aforementioned samples, many of which are only candidates themselves, to assess the contamination and purity of our method to the extent that is possible.
@@ 4.3 Data expectations
As a baseline for our morphology-based approach, it is first necessary to verify that the Euclid VIS images contain sufficient information to resolve the contrast between the compact central emission characteristic of an AGN and the broader light profiles of typical galaxies. We analysed the relationship between the brightest central pixel and its surrounding pixels to quantify a difference between normal galaxies and AGN. By producing ratios comparing the central pixel to 1 and 2-pixels-wide regions around the central pixel, we formed a rate of change (ROC) metric over sources, which is described by
@@
Pixels surrounding the brightest pixel, which share values similar to the maximum, will produce a ROC value closer to zero, implying that the source does not feature a prominent active component. Sources featuring a single very bright pixel surrounded by fainter pixels will have a ROC value closer to its maximum of 2. The produced distributions of sources, analysed according to both $I_{\scriptscriptstyle\rm E}$ magnitude and various selection criteria (Sect. 4.2), are shown in Fig. 3. As was expected, sources with a high-flux point spread function (PSF) light profile show the highest ROC values. This effect is seen in the star sample and the QSO-dominated selections (R90 WISE, Quaia, DESI-QSO, and Euclid-QSO). On the other hand, spectroscopically confirmed normal galaxies show very broad ROC distributions.
@@ 4.4 Addressing the use of asinh for image scaling
Figure 4: Distribution differences of the ROC value over each class. The narrow peaks from the non-galaxy class are significantly widened after applying the asinh transformation, causing a large overlap between classes.
Table 1: Percentage of galaxies that have a ROC value higher than the mean ROC of the other classes.
@@
AGN 0.92% 3.98%
QSO 0.13% 0.72%
Stars 0.12% 1.80%
Given its prevalence in galaxy morphology-based research (Lupton et al., 2004), the effect of using an $\operatorname{asinh}$ transformation on the data is also investigated. The function is often used to benefit visual inspections by reducing the difference between the brightest and faintest pixels. In our approach, however, when attempting to differentiate galaxy and non-galaxy sources, the $\operatorname{asinh}$ function becomes a hindrance, causing the distributions to more significantly overlap, as is shown in Fig. 4. To generate a measure for class overlap, we measured the ROC for each image, taking the mean of each non-galaxy class. We used the percentage of our true galaxy sample with a ROC value over the respective class mean as our overlap metric. The extent of the overlap is showcased in Table 1, where galaxy overlap increases by at least a factor of 4 and in the extreme case, nearly 15 times more galaxies overlap with stars.
In the following sections of the paper, we discuss the results of the diffusion model and analyse the performance of the inpainting pipeline that automatically learns to reconstruct the central emission expected for the normal galaxy population, allowing us to identify stars and AGN as outliers.
@@ 5 Results
Although the focus of the paper is on the model’s performance on small areas of a galaxy’s core, Repaint’s adaptability allows the pipeline to handle a variety of mask shapes and sizes. More general examples of the inpainting performance can be seen in Fig. 17. With larger masks, fewer pixels remain for conditioning, making the output more susceptible to artefacts. However, Lugmayr et al. (2022) demonstrate that this can be mitigated by increasing the number of resamples during the Repaint loop, at the cost of longer inference times.
@@ 5.1 Inpainting of AGN
For our inpainting, we used a $\times5$ pixel mask centred on the brightest pixel, following the constraints described in Sect. 3.1. Comparing the expected pixel values with the generated outputs, Fig. 5 shows the model’s performance. The assumption that the model will predict a fainter core than the original does appear to be true, with a near-global bias for underpredicting, as is shown by the dashed $y=x$ line. The majority of sources behave in a consistent way, with the exception of a smaller cluster of sources that predict a core more than an order of magnitude fainter than the original. This distinct cluster could provide sufficient candidates for AGN.
@@
To analyse the reconstruction error for each subset of class candidates, comparisons were made between the original pixels and inpainted reconstruction for each image. The MSE of the pixel values are shown in Fig. 6. Each subset was binned according to its $I_{\scriptscriptstyle\rm E}$ magnitude, providing a distinction between the relative brightness of each image.
@@
There is a clear distribution difference between the respective class candidates. Brighter sources, particularly stars and luminous QSOs, produce higher errors, followed by fainter QSOs and the lowest-magnitude AGN candidates. Galaxies and the faintest AGN candidates show significantly lower errors, showing the model’s tendancy to reconstruct sources without strong AGN features more effectively. This indicates that the model has learned the anticipated bias introduced by the uneven distribution within the dataset. While MSE is scale-dependent and influenced by the source’s VIS $I_{\scriptscriptstyle\rm E}$ magnitude, the separation between the errors from the galaxy class and those of the other classes suggests it remains useful as a selection metric. The presence of low-error AGN candidates may also indicate contamination within the AGN labels.
@@
It is acknowledged that because the model is trained on the full, unlabelled galaxy population, its prior distribution does contain information about AGN, even if they are rare. This could, in principle, lead the model to occasionally reconstruct AGN-like features, resulting in false negatives. However, the strong performance in separating known AGN and QSO populations, as is shown in Fig. 6, suggests that the prior is sufficiently dominated by non-active galaxies for this outlier detection framework to be effective.
@@
To showcase how the performance of the inpainting method impacts the overall image, Fig. 7 presents the input, output, and residuals for a selection of sources, including a close-up of the masked area both before and after inpainting. It is clear how bright central cores can be entirely removed with inpainting and replaced by regions consistent with the surrounding pixels. The model can also accurately recreate images that do not feature any significant core. This shows that the model has a good internal representation of galaxies and how they should behave.
@@
Figure 9: Threshold produced between the top right cluster of points and the largely galaxy-dominated cluster. The coloured dots represent the VIS $I_{\scriptscriptstyle\rm E}$ magnitude for each positive candidate for the respective selections. The grey dots show the values of all the images in the dataset. The majority of the sources within each non-galaxy class are captured with this simple linear boundary, allowing for high recall across all classes. The sources not captured are typically those of a higher VIS magnitude, and therefore fainter. The tail formed, where sources exhibit a maximum pixel difference ratio of $10^{0}$, shows outlier sources with brighter predictions after inpainting.
@@ 5.2 Max pixel difference ratio
Although the reconstruction error allows stars and a subset of AGN to stand out, a substantial overlap between suspected AGN and DESI-classified galaxies remains, especially at higher magnitudes. Even though measured on a pixel-by-pixel scale, MSE averages away specific pixel differences, leading to less nuanced differences between input and output. For example, if the reconstructed image ended up being the correct values but shifted by one pixel, this would be penalised heavily under MSE. From the perspective of this paper, such a scenario would represent a desirable outcome, as the model would correctly predict the presence of a bright, unresolved component, even if its exact location were shifted by a few pixels within the mask. Another example would be if only the brightest pixel were different in the reconstruction. The MSE would significantly reduce the impact of this measurement, potentially scoring similarly to a source that gave minor errors in every pixel within the mask.
@@
To ensure such cases are not overlooked, we measured the pixel value ratio between the brightest pixel from the input mask and the brightest pixel from the reconstruction. As is shown in Fig. 8, the produced distributions show a much tighter spread of values for many of the AGN selections, with a skew towards higher scores. The majority of the galaxy subset produces a ratio of less than 10, whereas a large population of suspected AGN from each selection features a larger ratio.
@@
Another important difference between this metric and MSE is how the results are relative, reducing the impact of magnitude on the recorded values. Many of the high-density peaks remain consistent, even as the magnitude of the sources increases. This implies that this metric is more robust at detecting fainter AGN than the MSE.
@@
The distribution also highlights the scale of the extreme values, with many sources showing pixel values approaching 100 times greater than their counterparts. Exploring these large disparities can provide insights into the types of sources that are significantly sensitive to the model’s capabilities. Sources producing the highest values for the ratio are those for which the reconstruction produces a cutout significantly fainter than the original image. As is seen in Fig. 20, this space is dominated by point-like objects. Having the sharp fall-off in pixel intensity in such a small area leads to the mask covering the majority of the source, if not the entirety. When the model attempts to reconstruct the missing pixels, it generates values that are coherent with the surrounding pixels. In these cases, the point-like source contributes very little to the surrounding pixels, leading to the generation of only background noise without any source.
@@
The behaviour of the previous examples matches the model’s expected behaviour and the paper’s initial assumptions. Applying inpainting to this task is expected to produce outputs that are similar to, or produce a flattened light profile of, the original pixels. However, a subset of sources exists that exhibit the opposite behaviour and generate a much brighter collection of pixels. Figure 21 shows examples of such sources. When analysing the sources that exhibit this behaviour, we found that they have overwhelmingly low-S/N images. Further exploration of the impact of S/N on the diffusion process is shown in Sect. 6.1.
@@
Table 2: Number of sources that produced brighter pixels in the masked area after repainting.
@@
The number of sources whose max pixel difference ratio is less than one for models trained on the two loss functions discussed in Sect. 2.3 is shown in Table 2. The normalised loss significantly reduces the number of sources exhibiting this behaviour. The normalised loss trained on the complete dataset was selected as our applied model because it further reduced the number of erroneous predictions.
@@ 5.3 Morphology-based performance
Given the variety of shapes and sizes of objects that will be witnessed using Euclid, it is important to verify that the model can handle the complexities of different galaxy morphologies. Utilising the refined morphology classifications provided from Euclid Collaboration: Walmsley et al. (2025), we provide examples of edge-on galaxies (disk-edge-on_yes_fraction >0.5>0.5), spirals (has-spiral-arms_yes_fraction >0.7>0.7), and mergers (merging_merger_fraction >0.3>0.3) in Figs. 22, 23, LABEL:, and 24, respectively.
@@ 5.4 Creating a classifier
Due to the difference in distributions between the two metrics, any non-overlapping cases can be found when used in combination. As is seen in Fig. 9, there is a cluster of sources in the top right. The key aspect of this cluster is its clear boundary from the DESI galaxy selection. This is vital for ensuring high reliability in classifications. We created a simple decision boundary between the two clusters of points, as is shown in Fig. 9. Although a more fine-tuned boundary would be possible, especially when incorporating a secondary machine-learning model, it is important to show the benefits of utilising the inpainting method through only its error metrics. The equation of our linear boundary is $y=5.8x^{-0.06}$.
@@ 5.5 Variance of reconstructions
As part of a single application of the pipeline, seven samples are generated. To ensure that any low variance in samples is not caused by the model simply ‘remembering’ the true image from its training examples, all resampling is applied on images the model has never seen before. The mean and variance over a normalised-by-brightness MSE are shown in Fig. 10.
The top figure shows that, irrespective of magnitude, the model achieves a similar relative median error across all images, highlighting the impact of the normalised loss throughout training. However, the difference in the scheduler over images of different quality and brightness is evident in the bottom figure. The consistency of generated outputs varies drastically as fainter (and therefore typically lower S/N) images produce widely different pixels when resampled.
Figure 10: Variations in the output error from repeat samples of the same fixed pixels. Over seven separate inpainting runs, the model shows a consistent median error when normalised by each respective magnitude (top panel). The variance in the output errors is less consistent with the brighter, lower magnitudes, which provide more robust inpainting than those of fainter images.
Figure 11: Comparing the PSF total flux with the Srsic flux provides a clear separation between the galaxy class and the majority of the non-galaxy selection. This allows for a simple linear boundary for classifying potential AGN, with the majority being point-like objects. However, it is most apparent with selections from AllWISE R90 and X-ray-selected AGN that many sources would not be selected using these features. The coloured dots represent the VIS $I_{\scriptscriptstyle\rm E}$ magnitude for each positive candidate for the respective selections. The grey dots show the scores of all the images in the dataset.